It is known that the height of ten‐year old boys is normally distributed with a population standard deviation of 4.24 cm. If a random sample of 64 ten‐year old boys are considered what is the probability that the average height of this sample of ten‐year old boys will differ by no more than 1.15 cm in either direction from the unknown population mean height of ten‐year old boys?
Answer #1
sample size       =n= 64
std error=σ=σ/√n= 0.5300

probability that the average height of this sample of ten‐year old boys will differ by no more than 1.15 cm in either direction :

=P(-1.15/0.53 <Z< 1.15/0.53)=P(-2.17 <Z<2.17)=0.9850-0.0150 =0.9700